• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Fansler

    Log entry time 01:07:36 on February 16, 2011

    Entry number 347101

    keyword=Cryo Call in

    The counting house called and said the left Q1 had dumped on liquid level and was not refilling. I went into the hall and checked all the vacuums and the liquid level meters. The left Dipole had an offset of -30%. I replaced the liquid level meter and the level jumped from 60% to 99%. We also found a problem with the PID loop on the LQ2. The valves were closing and allowing the inlet temps to reach 150 degrees. The temp spike in conjunction with the Dipole starving the left arm magnets caused the Q1 to trip. We are attempting to refill the left Q1 at this time.